Terri asked her gradfather: How old are youn, grandpa? He said, I've lived a quarter of my life as a boy, a sixth of my life aas a young man, half of my life as a middle aged man, all those years I spent in my native country. Now you must add the past six years I've spent as a new immigrant. How old is Terri's grandfather?
1/4 + 1/6+ 1/12 =
3/12 + 2/12 + 6/12 = 11/12 + 6 ?????
I don't think I am on the right track. Please help

Answer by checkley77(12844)

x/4+x/6+x/2+6=x
(3x+2x+6x)/12+6=x
11x/12=x-6
11x=12(x-6)
11x=12x-72
1x-12x=-72
-x=-72
x=72 years old is the grandfather.